Three vertices of a parallelogram ABCD are A(4, 0, 3), B (3, 4, – 2) and C (– 2, 0, 1). Find the coordinates of the fourth vertex.

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student
we know AB=OA-OB=(1,-4,5)
also AB=-CD in a parallelogram
OD=OC+CD=(-2,0,1)+(-4,0,-3)
hence 4thvertex OD=(-6,0,-2)
Thanks and regards
Shaik Aasif
askIITians faculty